Can someone help me solve the solution of the system by using the elimination method to find the solution?
x + 5y = 20
-x + 8 = 19

Answer by mananth(16946)

x+5y=20 .............1
-x+8y=19 .............2

multiply (1)by -8
Multiply (2) by 5
-8x-40y= -160
-5x+40y =95
Add the two equations
-13x=-65
/-13
x=5.00
plug value of x in (1)
x+5y=20
5+5y=20
5y=20-5
5y=15
y=3